C= 2.85n + 30.52
a. find and interpret the C-intercept for the line
b. find and interpret the n-intercept for the line
c. graph the line for n ranging from 0 thru 20
thank you so so much. I have another one I will submit separately. would I be able to get this and a few other problems by Sunday, 9/12 or Monday 9/13? I'm in a panic. I can't tell you how grateful I would be. I just need to pass this algebra course.



Answer by Alan3354(69443)

C= 2.85n + 30.52
a. find and interpret the C-intercept for the line
At the C-int, n=0 (the interpretation)
C = 30.52
----------
b. find and interpret the n-intercept for the line
At the n-int, C = 0
0 = 2.85n + 30.52
n = -30.52/2.85
n =~ -10.7088
---------------
c. graph the line for n ranging from 0 thru 20
---------------
The C-intercept is the point (0,30.52)
At n = 10, C = 2.85*20 + 30.52 = 87.52
A 2nd point is (20,87.52)
It's a straight line, 2 points is sufficient to plot it.